### v2

# Anisotropic Green Coordinates

Abstract.
We live in a world filled with anisotropy, a ubiquitous characteristic of both natural and engineered systems. In this study, we concentrate on space deformation and introduce anisotropic Green coordinates, which provide versatile effects for cage-based and variational deformations in both two and three dimensions. The anisotropic Green coordinates are derived from the anisotropic Laplacian equation $\nabla\cdot(\mathbf{A}\nabla u)=0$, where $\mathbf{A}$ is a symmetric positive definite matrix. This equation belongs to the class of constant-coefficient second-order elliptic equations, exhibiting properties analogous to the Laplacian equation but incorporating the matrix $\mathbf{A}$ to characterize anisotropic behavior. Based on this equation, we establish the boundary integral formulation, which is subsequently discretized to derive anisotropic Green coordinates defined on the vertices and normals of oriented simplicial cages. Our method satisfies basic properties such as linear reproduction and translation invariance, and possesses closed-form expressions for both 2D and 3D scenarios. We also give an intuitive geometric interpretation of the approach, demonstrating that our method generates a quasi-conformal mapping. Furthermore, we derive the gradients and Hessians of the deformation coordinates and employ the local-global optimization framework to facilitate variational shape deformation, enabling flexible shape manipulation while achieving as-rigid-as-possible shape deformation. Experimental results demonstrate that anisotropic Green coordinates offer versatile and diverse deformation options, providing artists with enhanced flexibility and introducing a novel perspective on spatial deformation.
Anisotropic deformation, Green coordinates, Cage-based deformation, Variational shape deformation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Shape modelingccs: Computing methodologies Geometric deformation
1. Introduction
Space deformation is a basic problem in geometric modeling. Instead of directly modifying the object surface, space deformation treats the surface as an embedding in the ambient space and deforms the space through deformation handles (Ben-Chen et al., 2009). Cage-based deformation is a typical example in which the embedded object deforms according to the manipulation of the cage (Ströter et al., 2024). For each point $\eta$ within the cage, a series of functions $\phi_{i}(\eta)$ is constructed for each cage vertex $\mathbf{v}_{i}$. These functions satisfy the partition of unity property, $\sum_{i=1}^{n}{\phi_{i}(\eta)}=1$, and the linear reproduction property $\eta=\sum_{i=1}^{n}{\phi_{i}(\eta)}\mathbf{v}_{i}$. The functions $\phi_{i}(\eta)$ are commonly called the Generalized Barycentric Coordinates (GBC) (Floater, 2015; Hormann and Sukumar, 2017).
GBC is usually interpolatory and expresses the point inside the cage as an affine sum of the cage vertices. Although it demonstrates good consistency with the target cage, it introduces obvious shearing artifacts, particularly when the cage undergoes affine transformations (Lipman et al., 2008). In practice, it is also desirable for the deformed object to maintain a certain degree of conformality with the original object. Based on the Green’s third identity of harmonic functions, Lipman et al. (2008) propose Green coordinates, which take into account both the cage vertices $\mathbf{v}_{i}$ and the cage normals $\mathbf{n}_{j}$. Green coordinates satisfy $\eta=\sum_{i=1}^{n}{\phi_{i}(\eta)}\mathbf{v}_{i}+\sum_{j=1}^{m}\psi_{j}(\eta)\mathbf{n}_{j}$ for $\eta$ inside the cage. These coordinates enable shape-preserving deformations, characterized by angle-preserving in 2D and quasi-conformality in 3D. However, conformality is a strict and absolute condition, which also restricts the deformation diversity of Green coordinates once the source and target cages are fixed. In practical applications and artistic design, there is often a demand for more flexible and versatile deformation styles, which can be achieved through adjusting various parameters. In the real world, materials such as bones, wood, and fibers exhibit anisotropic structures, prompting users to seek deformations that are more “flexible” or more “rigid” along specific directions.
Based on previous observations, we propose to incorporate the anisotropic property into the traditional Green coordinates. The Laplacian operator is the composition of divergence and gradient, i.e., $\Delta=\nabla\cdot\nabla$. Therefore, we can introduce anisotropy within the divergence operator and consider the anisotropic Laplacian equation $\nabla\cdot(\mathbf{A}\nabla u)=0$, where $\mathbf{A}$ is a positive-definite matrix. The matrix $\mathbf{A}$ introduces varying scales in different directions, thereby incorporating anisotropic properties in the formulation. Furthermore, the anisotropic Laplacian equation also gives rise to corresponding boundary integral formulation and leads to the anisotropic Green’s identity, which form the foundation for constructing the anisotropic Green coordinates. We show that anisotropic Green coordinates satisfy linear reproduction and translation invariance properties, and produce versatile deformation results with varying control matrices $\mathbf{A}$. Additionally, we provide closed-form expressions for both 2D and 3D cases and offer intuitive geometric interpretations of the method. We also relate the distortion of the deformation mapping to the condition number of $\mathbf{A}$, and mathematically demonstrate that anisotropic Green coordinates exhibit quasi-conformal properties. Through extensive experiments, we validate that anisotropic Green coordinates provide artists with a broader range of deformation options and enhance the diversity of deformations choices. Moreover, appropriate selection of the matrix $\mathbf{A}$ enables desired effects, such as tighter adherence to the deformed cage or reduced isotropic and area distortion.
To further demonstrate the broad applicability of the closed-form anisotropic Green coordinates, we also derive their gradients and Hessians and perform variational shape deformation (Ben-Chen et al., 2009), which minimizes energy functional of the deformation map. This approach enables flexible control through space constraints and supports the as-rigid-as-possible deformation strategy. The incorporation of anisotropy further extends the applicability of the variational shape deformation framework to a wider range of scenarios.
2. Related Work
2.1. Cage-based Deformation
Cage-based deformation is a fundamental topic in geometry processing, where users manipulate a coarse control structure to drive the deformation of complex objects within the enclosed space (Ströter et al., 2024). Two critical factors in this problem are the geometry of the cage structure and the deformation coordinates defined on the cage. The most commonly used cage structure is the oriented simplicial surface, such as closed polygonal chains in 2D and triangle meshes in 3D (Ju et al., 2005; Lipman et al., 2008; Ben-Chen et al., 2009; Weber et al., 2012; Yan and Schaefer, 2019; Thiery et al., 2024). It is the first-order approximation of co-dimensional one surfaces. Other methods utilize high-order surfaces, such as polynomial or Bézier surfaces, as the cage structure (Li et al., 2013; Michel and Thiery, 2023; Qin et al., 2024; Liu et al., 2024; Lin and Chen, 2024; Michel et al., 2025; Xiao and Chen, 2025; Wu et al., 2025). Such structures enable more precise and flexible control over curved and bending deformations.
While the cage structure can influence the flexibility, the deformation effect is predominantly determined by the specific deformation coordinates established on the cage. The most widely utilized methods are generalized barycentric coordinates (GBC), which formulate interpolatory coordinates as a partition of unity and can accurately represent an arbitrary point $\eta$ within the cage. As pioneering works, Floater (2003), Floater et al. (2005), and Ju et al. (2005) construct closed-form mean value coordinates by projecting the enclosing cage onto a sphere centered at $\eta$ and calculating the contribution of each cage vertex based on this spherical projection. There are also other techniques for constructing GBC, such as using the Laplacian equation (Joshi et al., 2007), Poisson integral formula (Li and Hu, 2013), probability theory (Hormann and Sukumar, 2008; Chang et al., 2023), or stochastic sampling and Monte Carlo integration (de Goes and Desbrun, 2024). Additionally, transfinite coordinates (Belyaev, 2006; Dyken and Floater, 2009; Chang and Hormann, 2025) extend the concept of traditional discrete barycentric coordinates to continuous boundaries, offering a general theoretical framework.
In addition to barycentric coordinates, some other deformation coordinates consider both the cage vertices and cage normals, and therefore achieve shape-preserving deformations. Shape preservation means that the deformed shape maintains a certain degree of conformality with the original shape and avoids large shearing artifacts. The typical methods are the Green coordinates (Lipman et al., 2008; Lipman and Levin, 2009; Thiery and Boubekeur, 2022; Wu et al., 2025; Xiao and Chen, 2025), which utilize Green’s third identity to establish the deformation formulation and naturally satisfy linear reproduction and transformation invariance. Other techniques involve Cauchy coordinates (Weber et al., 2009; Lin and Chen, 2024; Liu et al., 2025b), which is demonstrated to be geometrically equivalent to Green coordinates in the planar domain. However, the deformation effects of Green coordinates and Cauchy coordinates are fixed. Specifically, once the source and target cages are determined, the deformation is uniquely defined. To address this limitation, some approach solve the biharmonic Dirichlet problem, which incorporates gradient variations into the formulation (Weber et al., 2012; Thiery et al., 2024; Liu et al., 2025a). Alternatively, Somigliana coordinates  (Chen et al., 2023) generalize Green coordinates based on the Navier–Cauchy equation of linear elasticity. This approach enables physically plausible volume control while preserving similarity invariance. However, it lacks a closed-form solution and requires a CUDA-dependent implementation. Another option is to use variational deformation techniques to optimize the deformation map, as discussed in the following subsection.
2.2. Variational Shape Deformation
Variational methods are widely used in formulating partial differential equations for practical problems, such as the minimal surface equation (Pinkall and Polthier, 1993). They address the most fundamental problem in functional analysis of finding a function $f$ that minimizes the energy functional $E(f)$.  Ben-Chen et al. (2009) introduce the variational harmonic map in space deformation, which defines $E(f)$ based on Green coordinates, enabling deformation control with space constraints and as-rigid-as-possible properties. The optimization problem is solved through a local-global minimization strategy (Sorkine and Alexa, 2007; Liu et al., 2008), which requires computing the Jacobian and Hessian of Green coordinates. Subsequent research has extended this technique by incorporating various coordinate systems or employing higher-order cage structures (Weber et al., 2012; Thiery et al., 2024; Liu et al., 2025b; Michel et al., 2025; Wu et al., 2025). Additionally, Dodik et al. (2023) introduce variational barycentric coordinates, a neural field-based framework that optimizes generalized barycentric coordinates by minimizing their total variation.
2.3. Anisotropy in Geometry Processing
Anisotropy is a fundamental phenomenon in both natural systems and physical processes, playing an important role in simulating directionally dependent materials and generating adaptive computational domains. Research in this field encompasses theoretical foundations, numerical methodologies, and practical applications.
At the mathematical fundamental level, Owhadi and Zhang (2007) establish a metric-based upscaling framework for elliptic PDEs of the form $\mathcal{L}u=\nabla\cdot(\mathbf{A}(\mathbf{x})\nabla u(\mathbf{x}))$ and rigorously analyze its mathematical properties within the Sobolev space.  Kharevych et al. (2009) construct anisotropic elasticity tensors on coarse meshes through harmonic homogenization to preserve fine-scale directional properties. Kim et al. (2019) focus on anisotropic hyperelasticity and propose an anisotropic strain invariant for the inversion-safe energy model, complemented by a rehabilitation technique to address degenerate elements. Lin et al. (2024) derive the anisotropic as-rigid-as-possible energy for elastic materials and provide a robust formulation suitable for finite element simulations.
Anisotropy is also well introduced in mesh processing. Chen and Desbrun (2022) expand Green’s functions by leveraging the anisotropic characteristics of general linear elastic materials, which allows for real-time control of material deformation without volumetric discretization. Ren et al. (2018) propose generating planar anisotropic meshes for a given metric tensor and establishing a quasi-conformal mapping to transform the anisotropic metric into a Euclidean metric. Zhong et al. (2013) embed the anisotropic Riemannian metrics into higher-dimensional isotropic spaces, achieving efficient particle-based anisotropic meshing. Ma et al. (2025) derive the anisotropic Gauss formulation and utilize this boundary integral formulation for point cloud normal orientation and surface reconstruction. This work focuses on geometry topics that are distinct from our method and only considers the anisotropic equations in diagonal forms.
3. Anisotropic Green Coordinates
In this section, we present a detailed derivation of the anisotropic Green coordinates. We begin by introducing the boundary integral formulation of the anisotropic Laplacian equation. It is non-trivial that, despite the introduction of anisotropy, we can still use boundary integrals to determine the coordinates of any point within the cage. Subsequently, we discretize this identity to obtain anisotropic Green coordinates that exhibit linear reproduction and translation invariance. Then, we analyze their properties.
3.1. Anisotropic Laplacian Equation
Green’s third identity can be derived from the Laplacian equation $\Delta u=0$, which serves as the foundation for establishing Green coordinates. Since $\Delta=\nabla\cdot\nabla$, we can introduce anisotropy within the divergence operator. Our method starts with the anisotropic Laplacian equation in $\mathbb{R}^{d},d\geq 2$, which is given by:
(1) $\nabla\cdot(\mathbf{A}\nabla u(\mathbf{x}))=0,$
where $\mathbf{A}$ is a $d\times d$ positive-definite matrix with constant coefficients. This equation is a special case of the second-order elliptic operator $\mathcal{L}u=\nabla\cdot(\mathbf{A}(\mathbf{x})\nabla u(\mathbf{x}))+\mathbf{b}(\mathbf{x})\cdot\nabla u(\mathbf{x})+c(\mathbf{x})u(\mathbf{x})$, and shares many similar properties with the classical Laplacian operator. When $\mathbf{A}$ is not constant, properties of solutions are generally studied in Sobolev spaces (Evans and Lawrence, 1964), and obtaining explicit solutions is often challenging. Thus, we focus on the scenario where $\mathbf{A}$ is a constant-coefficient matrix, and both the convection-enhancing term $\mathbf{b}(\mathbf{x})$ and the first-order term $c(\mathbf{x})$ are set to zero. It has been mathematically established that every differential operator $\mathcal{L}$ with constant coefficients has a fundamental solution (Folland, 1995).
We first introduce the concept of the square root of a positive definite matrix. It is known that every positive definite matrix $\mathbf{A}$ can be diagonalized as $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{\top}$, where
(2) $\Lambda=\mathrm{diag}(\lambda_{1},\lambda_{2},...,\lambda_{d})$
is a $d$-dimensional diagonal matrix with $\lambda_{i}>0$ for i=1,2,,di=1,2,...,d, and $\mathbf{P}$ is an orthogonal matrix satisfying $\mathbf{P}^{-1}=\mathbf{P}^{\top}$. By expressing
(3) $\Lambda^{1/2}=\mathrm{diag}(\sqrt{\lambda_{1}},\sqrt{\lambda_{2}},...,\sqrt{\lambda_{d}}),$
we can define the square root of $\mathbf{A}$ as $\mathbf{A}^{1/2}=\mathbf{P}\Lambda^{1/2}\mathbf{P}^{\top}$. It is then straightforward to verify that $(\mathbf{A}^{1/2})^{2}=\mathbf{A}$. $\mathbf{A}^{1/2}$ is also symmetric and positive definite.
Next, we will present some theorems related to the anisotropic Laplacian equation. While we believe these properties have been investigated in prior mathematical research, as they are closely related to the generalized Green’s function (Chen and Kuo, 2005), we are not aware of specific literature in the fields of cage or variational shape deformation that comprehensively addresses these results. Therefore, we provide a detailed discussion of them in the following.
Theorem 1.
The fundamental solution of the anisotropic Laplacian Equation $\nabla_{\xi}\cdot(\mathbf{A}\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta))=\delta(\xi-\eta)$ is:
(4) $G_{\mathbf{A}}(\xi,\eta)=\begin{cases}\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\log\sqrt{(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)},&d=2,\\ -\frac{1}{(d-2)\omega_{d}\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{\frac{2-d}{2}}},&d\geq 3,\end{cases}$
where $\delta(\mathbf{x})$ is the Dirac delta distribution satisfying $\delta(\mathbf{x})=0$ for $\mathbf{x}\neq\mathbf{0}$ and $\int_{\mathbb{R}^{d}}{\delta(\mathbf{x})=1}$, and $\omega_{d}={2\pi}^{d/2}/\Gamma(d/2)$ represents the surface area of the unit sphere in $\mathbb{R}^{d}$.
Proof.
We begin with the variable substitutions $\mathbf{x}=\mathbf{A}^{-1/2}\xi$ and $\mathbf{y}=\mathbf{A}^{-1/2}\eta$. According to Lemmas A.1 in the supplementary material, we have
(5) $\begin{gathered}\nabla_{\xi}\cdot(\mathbf{A}\nabla_{\xi}u)=\nabla_{\xi}\cdot(\mathbf{A}\mathbf{A}^{-1/2}\nabla_{\mathbf{x}}u)=\nabla_{\xi}\cdot(\mathbf{A}^{1/2}\nabla_{\mathbf{x}}u)\\ =\nabla_{\mathbf{x}}\cdot(\mathbf{A}^{-1/2}\mathbf{A}^{1/2}\nabla_{\mathbf{x}}u)=\Delta_{\mathbf{x}}u.\end{gathered}$
Since the Dirac delta distribution is a generalized function that violates the $\mathcal{C}^{2}$ condition of Lemma A.1, we introduce Lemma A.2 to establish the same variable substitution formula within the framework of distribution theory. Furthermore, within distribution theory, equality holds in the almost everywhere sense; thus, functions may differ on a set of measure zero. Nevertheless, for the fundamental solution, it is customary to seek a function that is smooth everywhere except at the singular point.
In the following, we denote by $G(\mathbf{u})=G(\mathbf{x},\mathbf{y})$ the fundamental solution of the isotropic Laplacian equation $\Delta_{\mathbf{x}}G(\mathbf{x},\mathbf{y})=\delta(\mathbf{x}-\mathbf{y})$, where $\mathbf{u}=\mathbf{x}-\mathbf{y}$, and by $G_{\mathbf{A}}(\xi,\eta)$ the fundamental solution of the anisotropic Laplacian equation $\nabla_{\xi}\cdot(\mathbf{A}\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta))=\delta(\xi-\eta)$ with respect to $\mathbf{A}$. According to Eq. (5), we know that $G_{\mathbf{A}}$ satisfying
(6) $\Delta_{\mathbf{x}}G_{\mathbf{A}}(\mathbf{x},\mathbf{y})=\delta(\mathbf{A}^{1/2}(\mathbf{x}-\mathbf{y}))=\frac{1}{\sqrt{\det{\mathbf{A}}}}\delta(\mathbf{x}-\mathbf{y}).$
The last equality follows from Lemma A.3 together with the identity $\det{\mathbf{A}^{1/2}}=|\det{\Lambda^{1/2}|}=\sqrt{\det{\mathbf{A}}}$. Therefore, we conclude that $\sqrt{\det{\mathbf{A}}}\ G_{\mathbf{A}}(\mathbf{x}-\mathbf{y})\$is the fundamental solution $G(\mathbf{u})$ of $d$-dimensional isotropic Laplacian equation, which is given by
(7) $G(\mathbf{u})=\begin{cases}\frac{1}{2\pi}\log|\mathbf{u}|,&d=2,\\ -\frac{1}{(d-2)\omega_{d}}\frac{1}{|\mathbf{u}|^{d-2}},&d\geq 3.\end{cases}$
Substituting $\mathbf{u}=\mathbf{x}-\mathbf{y}=\mathbf{A}^{-1/2}(\xi-\eta)$ to $G(\mathbf{u})$ and we complete the proof. ∎
Moreover, the gradient $\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta)$ can be evaluated explicitly as:
(8) $\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta)=\begin{cases}\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{-1}\mathbf{A}^{-1}(\xi-\eta)},&d=2,\\ \frac{1}{\omega_{d}\sqrt{\det{\mathbf{A}}}}{{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]}^{-\frac{d}{2}}\mathbf{A}^{-1}(\xi-\eta)},&d\geq 3,\end{cases}$
will also be used frequently in the subsequent analysis.
3.2. Construction of Anisotropic Green Coordinates
Before deriving the anisotropic Green coordinates, we first revisit the construction of the classical isotropic Green coordinates, which are defined on an oriented simplicial surface cage $P=(\mathbb{V},\mathbb{T})$ (Lipman et al., 2008; Lipman and Levin, 2009), where $\mathbb{V}$ denotes the cage vertices and $\mathbb{T}$ represents the face normals. Green coordinates are constructed based on Green’s third identity. Specifically, when $\Omega$ is a bounded open set in $\mathbb{R}^{d}$ with $\mathcal{C}^{2}$ boundary, and $u(\eta)$ is a harmonic function on $\overline{\Omega}$, then for $\eta\in\Omega$, we have:
(9) $u(\eta)=\int_{\partial\Omega}{u(\xi)\frac{\partial G}{\partial\mathbf{n}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi}}-\int_{\partial\Omega}{G(\xi,\eta)\frac{\partial u}{\partial\mathbf{n}}(\xi)\ \mathrm{d}\sigma_{\xi}},$
where $G(\xi,\eta)$ is the fundamental solution of the isotropic Laplacian equation (7). The directional derivatives are taken along the outward normal.
Remark: The left hand side of Eq. (9) can be written as $c(\eta)u(\eta)$ if $\eta\in\mathbb{R}^{d}$ is not constrained in the open set $\Omega$ and $\partial\Omega$ is $\mathcal{C}^{2}$. The coefficient $c(\eta)$ satisfies $c(\eta)=1$ if $\eta\in\Omega$, $c(\eta)=1/2$ if $\eta\in\partial\Omega$, and $c(\eta)=0$ if $\eta\in\mathbb{R}^{d}\setminus\overline{\Omega}$. We ignore the boundary continuity condition for discrete inputs when $\eta\in\Omega$.
Figure 1. Illustration of the hat function in 2D and 3D.
Given the cage $P=(\mathbb{V},\mathbb{T})$, we can use above identity to reproduce any point $\eta$ inside the cage using the vertices $\{\mathbf{v}_{i}|i\in I_{\mathbb{V}}\}$ and outward normals $\{\mathbf{n}_{j}|j\in I_{\mathbb{T}}\}$ of the cage boundary as follows:
(10) $\eta=F(\eta,P)=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)\mathbf{v}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}(\eta)\mathbf{n}_{j}},\ \ \eta\in\Omega,$
where
(11) $\phi_{i}(\eta)$ $=\int_{\xi\in\mathcal{N}\{\mathbf{v}_{i}\}}{\Gamma_{\mathbf{v}_{i}}(\xi)\frac{\partial G}{\partial\mathbf{n}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi},i\in I_{\mathbb{V}}},$
(12) $\psi_{j}(\eta)$ $=-\int_{\xi\in t_{j}}{G(\xi,\eta)\ \mathrm{d}\sigma_{\xi},j\in I_{\mathbb{T}}}.$
In the above formulations, $I_{\mathbb{V}}$ and $I_{\mathbb{T}}$ represent the index set of the cage vertices and faces, respectively; $\mathcal{N}\{\mathbf{v}_{i}\}$ represents the faces adjacent to $\mathbf{v}_{i}$, and $\Gamma_{\mathbf{v}_{i}}(\xi)$ is a piecewise linear hat function, which takes the value one at $\mathbf{v}_{i}$, zero at all other vertices, and varies linearly across each face within $\mathcal{N}\{\mathbf{v}_{i}\}$. As illustrated in Fig. 1, the explicit form of $\Gamma_{\mathbf{v}_{i}}(\xi)$ in the 2D case is given by $\Gamma_{\mathbf{v}_{i}}(\xi)=\frac{\|\xi-\mathbf{v}_{i+1}\|}{\|\mathbf{v}_{i+1}-\mathbf{v}_{i}\|}$. In the 3D scenario, $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)$ is expressed as $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)=1-\frac{(\xi-\mathbf{v}_{j_{1}})\cdot\mathbf{h}_{j_{1}}}{\|\mathbf{h}_{j_{1}}\|^{2}}$, where $\mathbf{h}_{j_{1}}$ denotes the altitude corresponding to the edge opposite $\mathbf{v}_{j_{1}}$. Furthermore, $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)$ essentially represents the barycentric coordinate of $\xi$ relative to $\mathbf{v}_{j_{1}}$ within $\triangle\mathbf{v}_{j_{1}}\mathbf{v}_{j_{2}}\mathbf{v}_{j_{3}}$. That is, $\xi=\Gamma_{\mathbf{v}_{j_{1}}}(\xi)\mathbf{v}_{j_{1}}+\Gamma_{\mathbf{v}_{j_{2}}}(\xi)\mathbf{v}_{j_{2}}+\Gamma_{\mathbf{v}_{j_{3}}}(\xi)\mathbf{v}_{j_{3}}$ and $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)+\Gamma_{\mathbf{v}_{j_{2}}}(\xi)+\Gamma_{\mathbf{v}_{j_{3}}}(\xi)=1.0$.
When the source cage $P$ is deformed to the target cage $\tilde{P}$, the new location $\tilde{\eta}$ of $\eta$ is computed as:
(13) $\tilde{\eta}=F(\eta,\tilde{P})=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)\tilde{\mathbf{v}}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}(\eta)s_{j}\tilde{\mathbf{n}}_{j}},$
where $\tilde{\mathbf{v}}_{i}$ and $\tilde{\mathbf{n}}_{j}$ denote the vertices and normals of the deformed cage, respectively. Throughout this work, the tilde symbol is used to represent quantities associated with the deformed value or target cage. The scale factor $s_{j}$ preserves scale invariance in 2D and quasiconformality in 3D, as detailed in Section 3 of  (Lipman et al., 2008).
Next, we derive the anisotropic boundary integral formulation of the anisotropic Laplacian equation, which establishes the mathematical foundation for our anisotropic Green coordinates.
Theorem 2.
If $\Omega$ is a bounded open set in $\mathbb{R}^{d}$ with $\mathcal{C}^{2}$ boundary, and $u(\xi)$ is a $\mathcal{C}^{2}$ scalar function satisfying $\nabla\cdot(\mathbf{A}\nabla u(\mathbf{\xi}))=0$ on $\overline{\Omega}$, where $\mathbf{A}$ is symmetric positive definite. Then, for $\eta\in\Omega$, we have
(14) $\begin{gathered}u(\eta)=\int_{\partial\Omega}{u(\xi)\ (\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}}\\ -\int_{\partial\Omega}{G_{\mathbf{A}}(\xi,\eta)(\mathbf{A}\nabla u(\xi)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}},\end{gathered}$
where $G_{\mathbf{A}}(\xi,\eta)$ is the fundamental solution of the anisotropic Laplacian equation $\nabla_{\xi}\cdot(\mathbf{A}\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta))=\delta(\xi-\eta)$, and $\mathbf{n}(\xi)$ the outward normal of $\partial\Omega$ at $\xi$.
Proof.
We begin the proof by applying the divergence theorem. Let $\Omega$ be a bounded open set in $\mathbb{R}^{d}$ and $\mathbf{F}$ is a $\mathcal{C}^{2}$ vector-valued function defined on the closure $\overline{\Omega}$. Then,
(15) $\int_{\Omega}\nabla\cdot\mathbf{F}\ \mathrm{d}V=\int_{\partial\Omega}\mathbf{F}\cdot\mathbf{n}\ \mathrm{d}\sigma,$
where $\mathbf{n}$ is the normal vector of $\partial\Omega$. Let $\mathbf{F}=u(\mathbf{A}\nabla v)$, where $u,v$ are two scalar functions on $\overline{\Omega}$. By applying the rule for computing the divergence of function products, we obtain the following equation:
(16) $\nabla\cdot\mathbf{F}=\nabla u\cdot(\mathbf{A}\nabla v)+u\nabla\cdot(\mathbf{A}\nabla v).$
Combining Eqs. (15) and (16), we get
(17) $\int_{\Omega}{\nabla u\cdot(\mathbf{A}\nabla v)+u\nabla\cdot(\mathbf{A}\nabla v)}\ \mathrm{d}V=\int_{\partial\Omega}{u(\mathbf{A}\nabla v\cdot\mathbf{n})}\ \mathrm{d}\sigma.$
Exchanging $u$ and $v$ in the above equation and obtain
(18) $\int_{\Omega}{\nabla v\cdot(\mathbf{A}\nabla u)+v\nabla\cdot(\mathbf{A}\nabla u)}\ \mathrm{d}V=\int_{\partial\Omega}{v(\mathbf{A}\nabla u\cdot\mathbf{n})}\ \mathrm{d}\sigma.$
Then, we can subtract Eq. (18) from Eq. (17). Since $\mathbf{A}$ is symmetric,
(19) $\nabla u\cdot(\mathbf{A}\nabla v)={\nabla u}^{\top}\mathbf{A}\nabla v={\nabla u}^{\top}\mathbf{A}^{\top}\nabla v=(\mathbf{A}\nabla u)^{\top}\nabla v=\nabla v\cdot(\mathbf{A}\nabla u),$
Therefore, we have
(20) $\int_{\Omega}{u\nabla\cdot(\mathbf{A}\nabla v)-v\nabla\cdot(\mathbf{A}\nabla u)\ \mathrm{d}V}=\int_{\partial\Omega}{u(\mathbf{A}\nabla v)\cdot\mathbf{n}-v(\mathbf{A}\nabla u)\cdot\mathbf{n}}\ \mathrm{d}\sigma.$
Let $v(\xi)=G_{\mathbf{A}}(\xi,\eta)$, where $\eta\in\Omega$ is fixed, and the integral is taken with respect to the variable $\xi$. Then, $\nabla\cdot(\mathbf{A}\nabla v(\xi))=\delta(\xi-\eta)$, and we have
(21) $\int_{\Omega}{u(\xi)\ \nabla\cdot(\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta))}\ \mathrm{d}V_{\xi}=\int_{\Omega}{u(\xi)\delta(\xi-\eta)\ \mathrm{d}V_{\xi}}=u(\eta).$
Consequently, when $u$ satisfies $\nabla\cdot(\mathbf{A}\nabla u(\mathbf{\xi}))=0$, we obtain Equation (14) and complete the proof. ∎
Remark: From a strict mathematical perspective, the proof above remains not rigorous enough, as $G_{\mathbf{A}}$ is singular at $\xi=\eta$ when performing the volume integration. The standard approach is to excise $B_{\epsilon}(\eta)$ centered at $\eta$ and let $\varepsilon\rightarrow 0$, or to construct a sequence of functions that converges pointwise to the fundamental solution while remaining free of singularities (replacing $\|\xi-\eta\|$ with $(\|\xi-\eta\|^{2}+\varepsilon^{2})^{1/2}$. Here, however, we prove the result directly in the sense of distributions.
We can derive the anisotropic Green coordinates with the natural linear reproduction property by setting $u(\eta)=\eta$ in Eq. (14). There is a remark that $u$ is a scalar function in Eq. (14). However, we can apply the identity to each coordinate component separately. Specifically, let $u_{k}(\mathbf{\eta})=\eta_{k}$, which returns the $k$-th component of $\eta=(\eta_{1},\eta_{2},...,\eta_{d})^{\top}$. Moreover, denote $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d})$, where $\mathbf{a}_{k}$ is the $k$-th column of $\mathbf{A}$. Then, $\mathbf{A}\nabla u_{k}(\xi)\cdot\mathbf{n}(\xi)=(\mathbf{a}_{k})^{\top}\mathbf{n}(\xi)$. We can then combine the $d$ components to obtain $\eta=(u_{1}(\eta),...,u_{d}(\eta))^{\top}$. Moreover, the composed normal term on the right-hand side is $\mathbf{A}^{\top}\mathbf{n}_{j}=\mathbf{A}\mathbf{n}_{j}$. Considering the boundary integration region as the cage boundary, we have the following equations:
(22) $\eta=$ $\sum_{j\in I_{\mathbb{T}}}\Big({\int_{t_{j}}{\xi\ (\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}}-\int_{t_{j}}{G_{\mathbf{A}}(\xi,\eta)\ \mathbf{A}\mathbf{n}(\xi)\ \mathrm{d}\sigma_{\xi}}}\Big)$
$=$ $\sum_{j\in I_{\mathbb{T}}}\sum_{\mathbf{v}_{i}\in\mathbb{V}(t_{j})}{\mathbf{v}_{i}\Big(\int_{t_{j}}{\Gamma_{i}(\xi)\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi)\ \mathrm{d}\sigma_{\xi}}\Big)}$
$-\sum_{j\in I_{\mathbb{T}}}{\mathbf{A}\mathbf{n}_{j}}\Big(\int_{t_{j}}{G_{\mathbf{A}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi}}\Big),$
where $I_{\mathbb{T}}$ denotes the index set of the cage faces, and $\mathbb{V}(t_{j})$ represents the vertex set of $t_{j}$. By rearranging the above equation, the anisotropic Green coordinates can be expressed as:
(23) $\eta=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\mathbf{v}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)\ \mathbf{A}\mathbf{n}_{j}},\eta\in\Omega,$
where
(24) $\phi_{i}^{\mathbf{A}}(\eta)$ $=\int_{\xi\in\mathcal{N}\{\mathbf{v}_{i}\}}{\Gamma_{i}(\xi)\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi)\ \mathrm{d}\sigma_{\xi},\ i\in I_{\mathbb{V}}},$
(25) $\psi_{j}^{\mathbf{A}}(\eta)$ $=-\int_{\xi\in t_{j}}{G_{\mathbf{A}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi},\ j\in I_{\mathbb{T}}}.$
Compared to the traditional Green coordinates (Eqs (10) to (12)), the integral expressions of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are now related to $G_{\mathbf{A}}(\xi,\eta)$ instead of $G(\xi,\eta)$. Moreover, the Neumann term of Eq. (23), which is constructed based on the cage normals, now becomes $\mathbf{A}\mathbf{n}_{j}$. When the source cage is deformed to $\tilde{P}$ with new vertices $\tilde{\mathbf{v}}_{i}$ and face normals $\tilde{\mathbf{n}}_{j}$, the deformed position $\tilde{\eta}$ of $\eta$ is given by:
(26) $\tilde{\eta}=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\tilde{\mathbf{v}}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)s_{j}\mathbf{A}\tilde{\mathbf{n}}_{j}},$
where $s_{j}$ is the scale factor analogous to that in (Lipman et al., 2008). A detailed discussion of this term will be provided in the following subsection.
3.3. Properties
To analyze the deformation properties of the anisotropic Green coordinates, we first revisit the properties of the isotropic case, as examined in (Lipman et al., 2008; Lipman and Levin, 2009). Specifically, these properties include: (1) Linear reproduction: $\eta=F(P,\eta)$ for $\eta\in\Omega$; (2) Translation invariance: $\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)=1}$ for $\eta\in\Omega$; (3) Rotation and scale invariance: for an affine transformation $T$, $T\eta=F(\eta,TP)$; (4) Shape preservation: the deformation is angle-preserving in 2D and quasi-conformal in 3D; and (5) Smoothness: $\phi_{i}(\eta)$ and $\psi_{j}(\eta)$ are $C^{\infty}$ for $\eta\in\Omega$. We observe that the anisotropic Green coordinates retain the properties of linear reproduction, translation invariance, scale invariance, quasi-conformality and smoothness. However, it does not satisfy the rotation invariance. Additionally, in 2D, these coordinates are not angle-preserving but exhibit quasi-conformality.
Linear reproduction: The anisotropic boundary integral formula naturally satisfies the linear reproduction property. As shown in Eq. (26), when the source and target cages are identical and $s_{j}=1$, substituting the coordinate expressions back yields Eq. (22), verifying that $\tilde{\eta}=\eta$.
Translation invariance: Translation invariance follows from the fact that $\phi_{i}^{\mathbf{A}}(\eta)$ satisfies the partition of unity property, i.e., $\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)=1}$. This property can be proven by setting $u(\eta)=1$ in Eq. (14), which yields:
(27) $\int_{\partial\Omega}(\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}=1,\ \ \eta\in\Omega.$
Furthermore, $\Gamma_{\mathbf{v}_{i},t_{j}}(\eta)$ in the expression for $\phi_{i}^{\mathbf{A}}(\eta)$ is actually the barycentric coordinates of $\mathbf{v}_{i}$ with respect to $t_{j}$. Therefore,
(28) $\sum_{i\in I_{\mathbb{V}}}\phi_{i}^{\mathbf{A}}(\eta)=$ $\sum_{j\in I_{\mathbb{T}}}\sum_{\mathbf{v}_{i}\in\mathbb{V}(t_{j})}\Big({\int_{t_{j}}{\Gamma_{\mathbf{v}_{i},t_{j}}(\xi)\ (\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}\Big)}}$
$=$ $\int_{\partial\Omega}(\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}=1.$
[…]
(29) $s_{j}=\sqrt{\frac{\sigma_{1}^{2}+\sigma_{2}^{2}}{2}}=\frac{\sqrt{\|\tilde{u}\|^{2}\|v\|^{2}-2(\tilde{u}\cdot\tilde{v})+\|\tilde{v}\|^{2}\|u\|^{2}}}{\sqrt{8}\mathrm{area}(t_{j})},$
where $\sigma_{1}$ and $\sigma_{2}$ denote the singular values of the linear map that transforms $t_{j}$ to $\tilde{t}_{j}$ when restricted to a plane. Moreover, $u,v$ and $\tilde{u},\tilde{v}$ represent the edge vectors of $t_{j}$ and $\tilde{t}_{j}$, respectively.
Quasi-conformality Although anisotropic Green coordinates are not strictly conformal in two dimensions, their quasi-conformality in both 2D and 3D can be established through the analysis of their relationship with the isotropic formulation presented in Section 5. Consequently, we also include the analysis of quasi-conformality in Section 5.
Smoothness and generalized harmonicity We also have the property that $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are smooth for $\eta\in\Omega$. The reason is that $G_{\mathbf{A}}(\xi,\eta)$ is an elementary function with respect to $\eta$ and has no singularities within the integration domain as $\xi\neq\eta$. Consequently, $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are also smooth with respect to $\eta$, as they are obtained through differentiation and integration of $G_{\mathbf{A}}(\xi,\eta)$ with respect to $\xi$, and $\Gamma_{i}(\xi)$ is independent of $\eta$. Moreover, we also conclude that $\nabla_{\eta}\cdot(\mathbf{A}\nabla_{\eta}\phi_{i}^{\mathbf{A}}(\eta))=0$ and $\nabla_{\eta}\cdot(\mathbf{A}\nabla_{\eta}\psi_{j}^{\mathbf{A}}(\eta))=0$. We call this property generalized harmonic with respect to $\mathbf{A}$. This can be briefly proven as follows: Due to the smoothness of $G_{\mathbf{A}}(\xi,\eta)$ when $\xi\neq\eta$, differentiation with respect to $\eta$ can be interchanged with the integration with respect to $\xi$. Moreover, the gradient of $G_{\mathbf{A}}(\xi,\eta)$ with respect to $\eta$ is simply the negative of its gradient with respect to $\xi$. Given that $G_{\mathbf{A}}(\xi,\eta)$ is generalized harmonic with respect to $\mathbf{A}$ in $\xi$, it is also generalized harmonic with respect to $\eta$ in $\Omega$. It follows that $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ also inherit this property.
4. Derivation of Closed-form Anisotropic Green Coordinates
We have derived the deformation formula for anisotropic Green coordinates and investigated its fundamental properties. In this section, we present closed-form expressions for $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ in both 2D and 3D. These analytical solutions are more accurate and computationally efficient compared to numerical methods for evaluating the integrals. In the following derivations, we frequently use formulations for the substitution of area elements on curves and surfaces embedded in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. These formulations are detailed in Lemmas A.4, A.5, and A.6 of the supplementary material. While these results do not rely on advanced mathematical concepts, they are non-trivial and not immediately obvious. Additionally, Appendix B provides a detailed derivation of the closed-form expressions for classical isotropic Green coordinates and their gradients and Hessians, which serve as the foundation for the anisotropic formulation.
4.1. Closed-form Expressions for 2D Scenario
We begin by presenting the method in 2D. We assume that the 2D cage is a closed and oriented 1-mainfold in $\mathbb{R}^{2}$ with vertices $I_{\mathbb{V}}=\{\mathbf{v_{i}}\}_{i=1}^{N}$ and faces $I_{\mathbb{T}}=\{f_{j}|f_{j}=\overrightarrow{\mathbf{v}_{j}\mathbf{v}_{j+1}}\}$ arranged in counter-clockwise order. We denote $\mathbf{a}_{j}=\mathbf{v}_{j+1}-\mathbf{v}_{j}$. Then, the unit outward normal $\mathbf{n}_{j}$ of $f_{j}$ is given by $\mathbf{n}_{j}={\mathbf{a}_{j}}^{\bot}/\|{\mathbf{a}_{j}}^{\bot}\|$, where the superscript $\bot$” denotes the vector obtained by rotating the original vector 90 degrees clockwise. Then, $\psi_{j}^{\mathbf{A}}(\eta)$ has the following expression:
(30) $\psi_{j}^{\mathbf{A}}(\eta)=$ $-\int_{\xi\in f_{j}}G_{\mathbf{A}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi},$
(31) $=$ $\int_{\xi\in f_{j}}{-\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\log\sqrt{(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)}}\ \mathrm{d}\sigma_{\xi}.$
Figure 2. The illustration of a 2D cage and its mathematical notations.
The key step is to transform the calculation of the anisotropic formulation into the isotropic one through the variable substitution. Specifically, let $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j}=\mathbf{A}^{-1/2}\mathbf{v}_{j}$ and $\mathbf{v}^{\prime}_{j+1}=\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$. According to Lemma A.4 and Lemma A.5 of the supplementary material, we have $\mathrm{d}\sigma_{\mathbf{x}}=|\det\mathbf{A}^{-1/2}|\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\ \mathrm{d}\sigma_{\xi}$, where $\mathbf{n}_{j}$ is the unit normal of $f_{j}$. Denoting $f_{j}^{\prime}=\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$, we have
(32) $\psi_{j}^{\mathbf{A}}(\eta)=$ $\int_{\xi\in f_{j}}{-\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\log\sqrt{(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)}}\ \mathrm{d}\sigma_{\xi}$
$=$ $\frac{1}{|\det{\mathbf{A}^{-1/2}}|\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}\int_{\xi\in f_{j}^{\prime}}-{\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\log\|\mathbf{x}-\mathbf{y}\|}\ \mathrm{d}\sigma_{\mathbf{x}}$
$=$ $\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\int_{\xi\in f_{j}^{\prime}}{-\frac{1}{2\pi}\log\|\mathbf{x}-\mathbf{y}\|}\ \mathrm{d}\sigma_{\mathbf{x}}.$
[…]
(33) $Q(t)=$ $\left(t+\frac{\mathbf{a}_{j}\cdot\mathbf{b}_{j}}{\|\mathbf{a}_{j}\|^{2}}\right)\log\left(\|\mathbf{a}_{j}\|^{2}t^{2}+2(\mathbf{a}_{j}\cdot\mathbf{b}_{j})t+\|\mathbf{b}_{j}\|^{2}\right)$
$-\left(\frac{\|\mathbf{a}_{j}\|^{2}+(\mathbf{a}_{j}\cdot\mathbf{b}_{j})}{\|\mathbf{a}_{j}\|^{2}\sqrt{\|\mathbf{a}_{j}\|^{2}\|\mathbf{b}_{j}\|^{2}-(\mathbf{a}_{j}\cdot\mathbf{b}_{j})^{2}}}\right)\arctan\left(\frac{\|\mathbf{a}_{j}\|^{2}t+(\mathbf{a}_{j}\cdot\mathbf{b}_{j})}{\sqrt{\|\mathbf{a}_{j}\|^{2}\|\mathbf{b}_{j}\|^{2}-(\mathbf{a}_{j}\cdot\mathbf{b}_{j})^{2}}}\right).$
Please refer to Appendix B.1 for more details on this formula.
We now proceed to the Dirichlet term $\phi_{\mathbf{v}_{j}}^{\mathbf{A}}(\eta)$, which is obtained by summing the contributions over the edges associated with $\mathbf{v}_{i}$. We focus on the integral over a single edge $f_{j}$, denoted as $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)$. Recall that $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j}=\mathbf{A}^{-1/2}\mathbf{v}_{j}$, $\mathbf{v}^{\prime}_{j+1}=\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$, $f_{j}=\overrightarrow{\mathbf{v}_{j}\mathbf{v}_{j+1}}$ and $f_{j}^{\prime}=\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$. Let $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi)$ be the hat-function of $\xi$ with respect to $\mathbf{v}_{j}$ in $f_{j}$, and $\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x})$ be the hat-function of $\mathbf{x}$ with respect to $\mathbf{v}_{j}^{\prime}$ in $f_{j}^{\prime}$. We first show that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ for $\xi_{0}\in f_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Recall that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})$ is the barycentric coordinates of $\xi_{0}$ with respect to $\mathbf{v}_{j}$ in $f_{j}$, we obtain
(34) $\xi_{0}=\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})\mathbf{v}_{j}+\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi)\mathbf{v}_{j+1}$
and
(35) $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})+\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})=1.$
Then,
(36) $\mathbf{x}_{0}=$ $\mathbf{A}^{-1/2}\xi=\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})\mathbf{A}^{-1/2}\mathbf{v}_{j}+\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$
$=$ $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})\mathbf{v}_{j}^{\prime}+\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})\mathbf{v}_{j+1}^{\prime}.$
Therefore, $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0}),\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})$ and $\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0}),\Gamma_{\mathbf{v}_{j+1}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ are both barycentric coordinates of $\mathbf{x}_{0}$ in $f_{j}^{\prime}$. As established in Lemma A.7, the barycentric coordinates of a simplex are uniquely determined due to the affine independence of its vertices. This property distinguishes them from barycentric coordinates in arbitrary polygonal domains, where uniqueness is generally not guaranteed. Therefore, we conclude that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ for any $\xi_{0}\in f_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Moreover, we have $\mathrm{d}\sigma_{\xi}=|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{j}^{\prime}\|\mathrm{d}\sigma_{\mathbf{x}}$, where $\mathbf{n}_{j}^{\prime}$ is the unit normal of $\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$. Then, we can derive the closed-from solution of $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)$ as follows:
(37) $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)=$ $\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\int_{\xi\in f_{j}}{\Gamma_{\mathbf{v}_{j},f_{j}}(\xi)\frac{(\xi-\eta)\cdot\mathbf{n}_{j}}{(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)}}\mathrm{d}\sigma_{\xi}$
$=$ $\frac{|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{j}^{\prime}\|}{2\pi\sqrt{\det{\mathbf{A}}}}\int_{\mathbf{x}\in f_{j}^{\prime}}{\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x})\frac{\mathbf{A}^{1/2}(\mathbf{x}-\mathbf{y})\cdot(\frac{\mathbf{A}^{-1/2}\mathbf{n}_{j}^{\prime}}{\|\mathbf{A}^{-1/2}\mathbf{n}_{j}^{\prime}\|})}{(\mathbf{x}-\mathbf{y})^{\top}(\mathbf{x}-\mathbf{y})}}\ \mathrm{d}\mathbf{x}$
$=$ $\frac{1}{2\pi}\int_{\mathbf{x}\in f_{j}^{\prime}}{\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x})\frac{(\mathbf{x}-\mathbf{y})\cdot\mathbf{n}_{j}^{\prime}}{(\mathbf{x}-\mathbf{y})^{\top}(\mathbf{x}-\mathbf{y})}\ \mathrm{d}\mathbf{x}}.$
Denote $\mathbf{a}_{j}=\mathbf{v}_{j+1}^{\prime}-\mathbf{v}_{j}^{\prime}$ and $\mathbf{b}_{j}=\mathbf{v}_{j}^{\prime}-\mathbf{y}$. The term
(38) $\int_{\mathbf{x}\in f_{j}^{\prime}}{\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x})\frac{(\mathbf{x}-\mathbf{y})\cdot\mathbf{n}_{j}^{\prime}}{(\mathbf{x}-\mathbf{y})^{\top}(\mathbf{x}-\mathbf{y})}\ \mathrm{d}\mathbf{x}}$
is equal to $\frac{\|\mathbf{a}_{j}\|(\mathbf{b}_{j}\cdot\mathbf{n}_{j})}{2\pi}(W(1)-W(0))$, where
(39) $W(t)=$ $\frac{1}{2\|\mathbf{a}_{j}\|^{2}}\log(\|\mathbf{a}_{j}\|^{2}t^{2}+2t(\mathbf{a}_{j}\cdot\mathbf{b}_{j})+\|\mathbf{b}_{j}\|^{2})$
- $\left(\frac{\|\mathbf{a}_{j}\|^{2}+(\mathbf{a}_{j}\cdot\mathbf{b}_{j})}{\|\mathbf{a}_{j}\|^{2}\sqrt{\|\mathbf{a}_{j}\|^{2}\|\mathbf{b}_{j}\|^{2}-(\mathbf{a}_{j}\cdot\mathbf{b}_{j})^{2}}}\right)\arctan\left(\frac{\|\mathbf{a}_{j}\|^{2}t+(\mathbf{a}_{j}\cdot\mathbf{b}_{j})}{\sqrt{\|\mathbf{a}_{j}\|^{2}\|\mathbf{b}_{j}\|^{2}-(\mathbf{a}_{j}\cdot\mathbf{b}_{j})^{2}}}\right).$
4.2. Closed-form Expressions for 3D Scenario
We can employ a similar methodology to derive the closed-form expression for the 3D case. Figure 3 depicts a tetrahedron formed by a triangular face $t_{j}=\triangle\mathbf{v}_{j_{1}}\mathbf{v}_{j_{2}}\mathbf{v}_{j_{3}}$ of the cage and an interior point $\eta$ within the cage. Let $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j_{1}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{1}}$, $\mathbf{v}^{\prime}_{j_{2}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{2}}$ and $\mathbf{v}^{\prime}_{j_{3}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{3}}$. According to Lemma A.6, we have $\mathrm{d}\sigma_{\mathbf{x}}=|\det\mathbf{A}^{-1/2}|\|\mathbf{A}^{1/2}\mathbf{n}_{t}\|\ \mathrm{d}\sigma_{\xi}$, where $\mathbf{n}_{t}$ denotes the unit outward normal vector of $t_{j}$. Furthermore, denoting $t_{j}^{\prime}=\triangle\mathbf{v}_{j_{1}}^{\prime}\mathbf{v}_{j_{2}}^{\prime}\mathbf{v}_{j_{3}}^{\prime}$, and its normal as $\mathbf{n}_{t}^{\prime}$. Then, the Neumann term $\psi_{j}^{\mathbf{A}}(\eta)$ on $t_{j}$ can be computed as follows:
(40) $\psi_{j}^{\mathbf{A}}(\eta)=$ $\int_{\xi\in t_{j}}{\frac{1}{4\pi\sqrt{\det{\mathbf{A}}}}\frac{1}{(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)}}\ \mathrm{d}\sigma_{\xi}$
$=$ $\frac{1}{|\det{\mathbf{A}^{-1/2}}|\|\mathbf{A}^{1/2}\mathbf{n}_{t}\|}\int_{\mathbf{x}\in t_{j}^{\prime}}{\frac{1}{4\pi\sqrt{\det{\mathbf{A}}}}\frac{1}{\|\mathbf{x}-\mathbf{y}\|}}\ \mathrm{d}\sigma_{\mathbf{x}}$
$=$ $\frac{1}{\sqrt{\mathbf{n}_{t}^{\top}\mathbf{A}\mathbf{n}_{t}}}\int_{\mathbf{x}\in t_{j}^{\prime}}{\frac{1}{4\pi}\frac{1}{\|\mathbf{x}-\mathbf{y}\|}}\ \mathrm{d}\sigma_{\mathbf{x}}.$
Figure 3. A tetrahedron formed by a triangular face $t_{j}=\triangle\mathbf{v}_{j1}\mathbf{v}_{j2}\mathbf{v}_{j3}$ of a cage and the query point $\eta$.
For the Dirichlet term, we denote by $\phi_{\mathbf{v}_{j_{1}},t_{j}}^{\mathbf{A}}(\eta)$ the contribution of $t_{j}$ to $\phi_{\mathbf{v}_{j_{1}}}^{\mathbf{A}}(\eta)$. According to Lemma A.6, we have $\mathrm{d}\sigma_{\xi}=|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{t}^{\prime}\|\mathrm{d}\sigma_{\mathbf{x}}$. Based on the uniqueness of the barycentric coordinates of a simplex (Lemma A.7), it holds that $\Gamma_{\mathbf{v}_{j_{1}},t_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j_{1}}^{\prime},t_{j}^{\prime}}(\mathbf{x}_{0})$ for $\xi_{0}\in t_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Therefore,
(41) $\phi_{\mathbf{v}_{j_{1}},t_{j}}^{\mathbf{A}}(\eta)=$ $\frac{1}{4\pi\sqrt{\det{\mathbf{A}}}}\int_{\xi\in t_{j}}{\Gamma_{\mathbf{v}_{j_{1}},t_{j}}(\xi)\frac{(\xi-\eta)\cdot\mathbf{n}_{t}}{[(\xi-\eta)^{\top}\mathbf{A}^{-1}(\xi-\eta)]^{\frac{3}{2}}}}\mathrm{d}\sigma_{\xi}$
$=$ $\frac{|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{t}^{\prime}\|}{4\pi\sqrt{\det{\mathbf{A}}}}\int_{\mathbf{x}\in t_{j}^{\prime}}{\Gamma_{\mathbf{v}_{j_{1}}^{\prime},t_{j}^{\prime}}(\mathbf{x})\frac{\mathbf{A}^{1/2}(\mathbf{x}-\mathbf{y})\cdot(\frac{\mathbf{A}^{-1/2}\mathbf{n}_{t}^{\prime}}{\|\mathbf{A}^{-1/2}\mathbf{n}_{t}^{\prime}\|})}{[(\mathbf{x}-\mathbf{y})^{\top}(\mathbf{x}-\mathbf{y})]^{\frac{3}{2}}}}\ \mathrm{d}\mathbf{x}$
$=$ $\frac{1}{4\pi}\int_{\mathbf{x}\in t_{j}^{\prime}}{\Gamma_{\mathbf{v}_{j_{1}}^{\prime},t_{j}^{\prime}}(\mathbf{x})\frac{(\mathbf{x}-\mathbf{y})\cdot\mathbf{n}_{t}^{\prime}}{[(\mathbf{x}-\mathbf{y})^{\top}(\mathbf{x}-\mathbf{y})]^{\frac{3}{2}}}\ \mathrm{d}\mathbf{x}}.$
Then, we transform the computation of the anisotropic formulation into an isotropic one. The calculation of 3D isotropic Green coordinates is somewhat complex, and we provide the details in Appendix B.2 of the supplementary material. These derivations are primarily based on the findings of  (Ben-Chen et al., 2009) and  (Urago, 2000), but we provide a more comprehensive derivation.
5. Discussion
In this section, we mainly derive the geometric interpretation of anisotropic Green coordinates in comparison to the isotropic counterparts, and demonstrate the quasi-conformality of our method. Furthermore, we emphasize the essential value of addressing problems from an anisotropic perspective.
We have the following observation when ignoring the effect of the scale factor $s_{j}$: the two approaches below are equivalent:
  1. (1)
    Given an input, first apply a global multiplication by $\mathbf{A}^{-1/2}$ to the source cage, target cage, and the object; subsequently deform the object using the isotropic Green coordinates; and finally multiply globally by $\mathbf{A}^{1/2}$.
  2. (2)
    Employ the anisotropic Green coordinates for the source and target cages.
We will present the derivation of this equivalence subsequently. Despite the equivalence, we believe that there are still important points to illustrate the value of deriving the anisotropic Green coordinates. First, our method provides a geometric interpretation of Green’s identity derived from the anisotropic Laplace equation. Second, as deformation coordinates, they need to be expressed as functions relative to the original cage, rather than the cage has undergone a global transformation. Moreover, when performing the variational method for deformation control (which will be presented in Section 6), it is more intuitive to analyze directly from the closed-form anisotropic formulation as well as its gradient and Hessian. Most importantly, during user interaction, the deformation coordinates $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ typically need to be calculated only once for the input source cage, and subsequent deformations do not require recalculation. If calculated directly using anisotropic Green coordinates, according to Eq. (26), only $s_{j}$, $\mathbf{A}\tilde{\mathbf{n}}_{j}$ and $\tilde{\eta}$ are required to be computed every time. In contrast, if calculated from the perspective of the global transform, it is necessary to first compute the transformed target cage and normals, which increases the computational overhead. Furthermore, this study introduces a new perspective on cage-based and variational deformation. If the matrix $\mathbf{A}$ can be optimized, there is potential to further enhance the practical value of the method, which remains a direction for our future work.
Next, we refer to the deformation effect obtained by “globally multiplying by $\mathbf{A}^{-1/2}$, performing deformation using isotropic Green coordinates, and then globally multiplying by $\mathbf{A}^{1/2}$”, as similarity Green deformation. The name stems from the fact that this process closely resembles a matrix similarity transformation. We denote the vertices and normals of the cage as $\{\mathbf{v}_{i}\}_{i=1}^{N}$ and $\{\mathbf{n}_{j}\}_{j=1}^{M}$. Moreover, we use the prime operator to denote the globally transformed values, e.g., $\mathbf{v}_{i}^{\prime}=\mathbf{A}^{-1/2}\mathbf{v}_{i}$, $\mathbf{n}_{j}^{\prime}=\frac{\mathbf{A}^{1/2}\mathbf{n}_{j}}{\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}$. Then, we can express the globally transformed values in terms of isotropic Green coordinates as:
(42) $\eta^{\prime}=\sum_{i\in I_{\mathbb{V}^{\prime}}}{\phi_{i}^{\prime}(\eta^{\prime})\mathbf{v}}_{i}^{\prime}+\sum_{j\in I_{\mathbb{T}^{\prime}}}{\psi_{j}^{\prime}(\eta^{\prime})\mathbf{n}_{j}^{\prime}}.$
The derivation of section 4.2 (Eqs. (32) and (37)) shows that $\phi_{i}^{\prime}(\eta^{\prime})=\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\prime}(\eta^{\prime})=\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\psi_{j}^{\mathbf{A}}(\eta)$, where $\mathbf{n}_{j}$ is the normal of $f_{j}$. Moreover, we have $\eta=\mathbf{A}^{1/2}\eta^{\prime}$. Then,
(43) $\eta$ $=\mathbf{A}^{1/2}\eta^{\prime}=\mathbf{A}^{1/2}(\sum_{i\in I_{\mathbb{V}^{\prime}}}{\phi_{i}^{\prime}(\eta^{\prime})\mathbf{v}}_{i}^{\prime}+\sum_{j\in I_{\mathbb{T}^{\prime}}}{\psi_{j}^{\prime}(\eta^{\prime})\mathbf{n}_{j}^{\prime}})$
$=\mathbf{A}^{1/2}(\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\mathbf{A}^{-1/2}\mathbf{v}}_{i}+\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)\frac{\mathbf{A}^{1/2}\mathbf{n}_{j}}{\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}})$
$=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\mathbf{v}}_{i}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)\mathbf{A}\mathbf{n}_{j}},$
and we show the equivalence between the anisotropic Green deformation and the similarity Green deformation.
Moreover, we can use the above analysis to demonstrate the quasi-conformality of our method. We denote the anisotropic deformation mapping as $f_{\mathbf{A}}(\eta)$, and the corresponding isotropic mapping as $f^{\prime}(\eta)$. Then,
(44) $f_{\mathbf{A}}(\eta)=\mathbf{A}^{1/2}f^{\prime}(\mathbf{A}^{-1/2}\eta).$
The Jacobian of $f_{\mathbf{A}}(\eta)$ is given by:
(45) $\mathbf{J}_{f_{\mathbf{A}}}(\eta)=\mathbf{A}^{1/2}\mathbf{J}_{f^{\prime}}(\mathbf{A}^{-1/2}\eta)\mathbf{A}^{-1/2},$
where $\mathbf{J}_{f^{\prime}}$ is the Jacobian of the isotropic formulation. Let $K_{\mathbf{A}}(\eta)$ denote the ratio of the maximum to the minimum singular values of the anisotropic deformation mapping, and $K^{\prime}(\eta)$ the corresponding isotropic ratio. Then,
(46) $K_{\mathbf{A}}(\eta)=$ $\frac{\sigma_{max}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))}{\sigma_{min}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))}=\text{cond}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))$
$=\text{cond}(\mathbf{A}^{1/2}\mathbf{J}_{f^{\prime}}(\mathbf{A}^{-1/2}\eta)\mathbf{A}^{-1/2})$
$\leq\text{cond}(\mathbf{A}^{1/2})\text{cond}(\mathbf{J}_{f^{\prime}}(\mathbf{A}^{-1/2}\eta))\text{cond}(\mathbf{A}^{-1/2})$
$=\text{cond}(\mathbf{A})K^{\prime}(\mathbf{A}^{-1/2}\eta),$
where $\sigma_{max}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))$ and $\sigma_{min}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))$ represent the maximum and minimum singular values of $\mathbf{J}_{f_{\mathbf{A}}}(\eta)$, and $\text{cond}(\mathbf{A})$ represents the condition number of $\mathbf{A}$. Therefore, when Green coordinates are conformal in 2D, anisotropic Green coordinates are quasi-conformal with $K_{\mathbf{A}}(\eta)\leq\text{cond}(\mathbf{A})$. When Green coordinates are quasi-conformal in 3D with a distortion bound $M$, anisotropic Green coordinates are quasi-conformal in 3D with a bound $\text{cond}(\mathbf{A})\cdot M$. Moreover, since $\mathbf{A}=\mathbf{P}^{\top}\Lambda\mathbf{P}$, the condition number $\text{cond}(\mathbf{A})$ is precisely the ratio of the largest to the smallest eigenvalues of $\Lambda$.
6. Variational Shape Deformation of Anisotropic Green Coordinates
In this section, we derive the expressions for the derivatives of anisotropic Green coordinates, specifically the gradients and Hessians of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$. These derivatives form the foundation for deformation control in the variational shape deformation framework introduced in Ben-Chen et al. (2009). In this method, an energy functional is defined to determine the optimal deformation mapping that satisfies user-specified constraints while simultaneously balancing smoothness and as-rigid-as-possible transformations. Previous work also demonstrates that the introduction of anisotropy does not conflict with our desire for rigid transformations at specific positions (Lin et al., 2024).
For a simplicial surface cage with vertices $\mathbb{V}$ and faces $\mathbb{T}$, we can define the deformation map $f_{\mathbf{a},\mathbf{b}}(\eta)$ as a linear combination of basis functions $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ based on the anisotropic Green coordinates as
(47) $f_{\mathbf{a},\mathbf{b}}(\eta)=\sum_{v\in\mathbb{V}}\mathbf{a}_{v}\phi_{v}^{\mathbf{A}}(\eta)+\sum_{t\in\mathbb{T}}\mathbf{b}_{t}\psi_{t}^{\mathbf{A}}(\eta),$
where $\mathbf{b}_{t}=\mathbf{A}\mathbf{n}_{t}$. Here, $\mathbf{a}_{v}\in\mathbb{R}^{d}$ represents the piecewise-linear coefficients defined on the cage vertices, while $\mathbf{n}_{t}\in\mathbb{R}^{d}$ denotes the piecewise-constant coefficients defined on the cage faces. To enlarge the deformation space, $\mathbf{a}_{v}$ and $\mathbf{b}_{t}$ are treated as independent variables and optimized separately. Since $\mathbf{A}$ is invertible, we can directly optimize $\mathbf{b}_{t}$, as $\mathbf{n}_{t}$ can be easily computed through $\mathbf{n}_{t}=\mathbf{A}^{-1}\mathbf{b}_{t}$.
To perform variational shape deformation, it is necessary to compute the gradients and Hessians of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$. The core idea is to apply Lemma A.1 from the supplementary material to transform the formulation into an isotropic form. For the computation of the isotropic formulation, please refer to Appendix B.
Let $t_{j}$ be a face of the cage and $t_{j}^{\prime}$ its transformation via left-multiplication of vertices by $\mathbf{A}^{-1/2}$. For any point $\eta_{0}$ on $t_{j}$ with normal $\mathbf{n}_{j}$, let $\mathbf{y}_{0}=\mathbf{A}^{-1/2}\eta_{0}$ be the corresponding point in $t_{j}^{\prime}$. Moreover, let $\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})$ denote the Neumann term of $t_{j}$ for the anisotropic Green coordinates at $\eta_{0}$, while $\psi_{t_{j}^{\prime}}(\mathbf{y}_{0})$ represents the Neumann term of $t_{j}^{\prime}$ for the isotropic formulation. Based on the derivations in Section 4.2, we obtain:
(48) $\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\psi_{t_{j}^{\prime}}(\mathbf{y}_{0}),$
Therefore, according to Lemma A.1, the gradient $\nabla_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})$ and Hessian $\mathbf{H}_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})$ can be calculated as:
(49) $\begin{gathered}\nabla_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\mathbf{A}^{-1/2}\nabla_{\mathbf{y}}\psi_{t_{j}^{\prime}}(\mathbf{y}_{0}),\\ \mathbf{H}_{\eta}\psi_{t_{j}}^{\mathbf{A}}(\eta_{0})=\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\mathbf{A}^{-1/2}\mathbf{H}_{\mathbf{y}}\psi_{t_{j}^{\prime}}(\mathbf{y}_{0})\mathbf{A}^{-1/2}.\end{gathered}$
For the Dirichlet term, we denote $\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})$ as the contribution of the integration over $t_{j}$ to $\phi_{i}^{\mathbf{A}}(\eta)$. Based on the derivations in Section 4.2, we have:
(50) $\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})=\phi_{i,t_{j}^{\prime}}(\mathbf{y}_{0}).$
Then, the gradient $\nabla_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})$ and Hessian $\mathbf{H}_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta)$ can be calculated as:
(51) $\begin{gathered}\nabla_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})=\mathbf{A}^{-1/2}\nabla_{\mathbf{y}}\phi_{i,t_{j}^{\prime}}(\mathbf{y}_{0})\\ \ \mathbf{H}_{\eta}\phi_{i,t_{j}}^{\mathbf{A}}(\eta_{0})=\mathbf{A}^{-1/2}\mathbf{H}_{\mathbf{y}}\phi_{i,t_{j}^{\prime}}(\mathbf{y}_{0})\mathbf{A}^{-1/2}.\end{gathered}$
Thus, the computation of the derivatives of the anisotropic Green coordinates is transformed into the isotropic formulation, and can be computed by the method described in Appendix B.
We now formulate the objective function for variational deformation. We express the deformation mapping $f_{\mathbf{a},\mathbf{b}}(\eta)$ in matrix form as:
(52) $f_{\mathbf{a},\mathbf{b}}(\eta)_{1\times d}=\begin{pmatrix}\Phi_{1\times n}&\Psi_{1\times m}\end{pmatrix}\begin{pmatrix}\mathbf{a}_{n\times d}\\ \mathbf{b}_{m\times d}\end{pmatrix}$,
where $n=|\mathbb{V}|$ denotes the number of vertices, and $m=|\mathbb{T}|$ represents the number of faces. The vector $\Phi$ is a row vector composed of all Dirichlet terms $\phi_{v}^{\mathbf{A}}(\eta)$, and $\Psi$ is a row vector composed of all Neumann terms $\psi_{t}^{\mathbf{A}}(\eta)$.
The Jacobian matrix $\mathbf{J}_{f}(\eta)$ and the Hessian $\mathbf{H}_{f}(\eta)$ of the deformation mapping $f_{\mathbf{a},\mathbf{b}}(\eta)$ are also linear in $\mathbf{a}$ and $\mathbf{b}$, and are expressed as follows:
(53) $(\mathbf{J}_{f}(\eta))_{d\times d}^{\top}=\begin{pmatrix}(\mathbf{G}_{\Phi})_{d\times n}&(\mathbf{G}_{\Psi})_{d\times m}\end{pmatrix}\begin{pmatrix}\mathbf{a}_{n\times d}\\ \mathbf{b}_{m\times d}\end{pmatrix}$,
where each column of $\mathbf{G}_{\Phi}$ captures the gradients of $\phi_{v}^{\mathbf{A}}(\eta)$, and each column of $\mathbf{G}_{\Psi}$ records the gradients of $\psi_{t}^{\mathbf{A}}(\eta)$. Moreover,
(54) $(\mathbf{H}_{f}(\eta))_{[d(d+1)/2]\times d}=\begin{pmatrix}(\mathbf{H}_{\Phi})_{[d(d+1)/2]\times n}&(\mathbf{H}_{\Psi})_{[d(d+1)/2]\times m}\end{pmatrix}\begin{pmatrix}\mathbf{a}_{n\times d}\\ \mathbf{b}_{m\times d}\end{pmatrix}$,
where each column of $\mathbf{H}_{\Phi}$ records the elements of the Hessian matrix of $\phi_{v}^{\mathbf{A}}(\eta)$, comprising $d(d+1)/2$ values. This vector is derived by flattening the symmetric Hessian matrix, retaining the diagonal elements and the upper-right components. Similarly, $\mathbf{H}_{\Psi}$ contains the elements of the Hessian matrix of $\psi_{t}^{\mathbf{A}}(\eta)$.
Based on the above formulation, we can now define an optimization framework. The energy functional $E(f_{\mathbf{a},\mathbf{b}})$ is designed to enforce as-rigid-as-possible (ARAP) behavior at specific points (typically medial axis points), ensure global smoothness (by minimizing the Hessian magnitude), and satisfy user constraints at designated points (i.e., deforming specific points to the user-defined positions). For discrete sampling, this problem is formulated as:
(55) $\begin{gathered}\min_{\mathbf{a},\mathbf{b},\mathbf{R}_{i}}E(f_{\mathbf{a},\mathbf{b}})=\sum_{i=1}^{d}\left\|\mathbf{J}_{f}(\mathbf{m}_{i})-\mathbf{R}_{i}\right\|_{F}^{2}+\lambda_{1}\sum_{i=1}^{r}\|f_{\mathbf{a},\mathbf{b}}(\mathbf{q}_{i})-\mathbf{f}_{i}\|_{2}^{2}\\ +\lambda_{2}\sum_{i=1}^{k}\left\|\mathbf{H}_{f}(\mathbf{w}_{i})\right\|_{F}^{2}+\lambda_{3}(\|\mathbf{a}-\mathbf{a}_{0}\|_{F}^{2}+\|\mathbf{b}-\mathbf{b}_{0}\|_{F}^{2}),\\ \mathrm{s.t.},\mathbf{R}_{i}^{\top}\mathbf{R}_{i}=\mathbf{I},\forall i=1,\dots,d.\end{gathered}$
Here, $\mathbf{a}_{0}$ and $\mathbf{b}_{0}$ denote the values of the input source cage. This formulation aligns more closely with (Michel et al., 2025) than with the earlier work (Ben-Chen et al., 2009), which employs soft positional constraints $\|f_{\mathbf{a},\mathbf{b}}(\mathbf{q}_{i})-\mathbf{f}_{i}\|_{2}^{2}$. Experiments demonstrate that soft constraints can effectively satisfy the requirements. Since the number of optimized variables is typically much larger than the number of constraints, we incorporate a regularization term $\|\mathbf{a}-\mathbf{a}_{0}\|_{F}^{2}+\|\mathbf{b}-\mathbf{b}_{0}\|_{F}^{2}$ to enhance the optimization stability. Further details regarding parameter selection and sampling point selection will be provided in the experiment section. We apply the local/global strategy to solve Eq. (55(Sorkine and Alexa, 2007; Liu et al., 2008; Ben-Chen et al., 2009). Specifically, in the local step, we fix $\mathbf{a}$ and $\mathbf{b}$ to solve for the optimal rotation matrices $\mathbf{R}_{i}$. This step formulates a well-known orthogonal Procrustes problem, which can be efficiently solved via singular value decomposition. In the global step, we fix $\mathbf{R}_{i}$ to solve for $\mathbf{a}$ and $\mathbf{b}$, resulting in an unconstrained least-squares problem that can be directly reduced to a linear system.
7. Experiments
This section presents the experimental results of our proposed method. We first focus on cage-based deformation and demonstrate the results obtained with different coefficient matrices $\mathbf{A}$. Subsequently, we showcase variational deformation outcomes subject to the same constraints using varying $\mathbf{A}$. All experiments are on a standard laptop equipped with an Intel CPU running at 2.40 GHz and 16GB of RAM.
We begin by describing how to obtain diverse matrices $\mathbf{A}$ for our experiments, focusing initially on the 2D scenario. By the spectral theorem, any symmetric positive definite matrix $\mathbf{A}$ can be decomposed into a diagonal matrix $\Lambda=\text{diag}(\lambda_{1},\lambda_{2})$ with $\lambda_{1},\lambda_{2}>0$ and an orthogonal matrix $\mathbf{P}$ such that $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{\top}$. Moreover, since flipping the sign of an eigenvector leaves $\mathbf{P}\Lambda\mathbf{P}^{\top}$ unchanged, we may choose $\mathbf{P}\in SO(2)$, i.e., a rotation matrix. Therefore, we parameterize $\mathbf{P}=\begin{pmatrix}\cos{\theta}&-\sin{\theta}\\ \sin{\theta}&\cos{\theta}\end{pmatrix}$ and denote $\mathbf{A}=\mathcal{A}(\theta,\lambda_{1},\lambda_{2})$. Moreover, since for any $c\in\mathbb{R}^{+}$, the equation $\nabla\cdot(\mathbf{A}\nabla u)=0$ is equivalent to $\nabla\cdot(c\mathbf{A}\nabla u)=0$ and leads to the same deformation mapping, we can normalize one of the diagonal entries of the matrix to $1.0$ in our formulation. Appendix C.1 provides some anisotropic matrices used in 2D deformation experiments together with the $\lambda_{1},\lambda_{2},\theta$ values.
The construction of 3D matrices follows a similar approach. The 3D diagonal matrix $\Lambda$ is defined by three positive parameters $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. The 3D rotation matrix $\mathbf{P}$ is parameterized by three Euler angles $\alpha,\beta,\gamma$ and can be expressed as $\mathbf{P}=\mathbf{R}_{z}(\alpha)\mathbf{R}_{y}(\beta)\mathbf{R}_{x}(\gamma)$, where
(56) $\mathbf{R}_{z}(\alpha)=\begin{pmatrix}\cos{\alpha}&-\sin{\alpha}&0\\ \sin{\alpha}&\cos{\alpha}&0\\ 0&0&1\end{pmatrix},\ \mathbf{R}_{y}(\beta)=\begin{pmatrix}\cos{\beta}&0&-\sin{\beta}\\ 0&1&0\\ \sin{\beta}&0&\cos{\beta}\end{pmatrix}$,
$\mathbf{R}_{x}(\gamma)=\begin{pmatrix}1&0&0\\ 0&\cos{\gamma}&-\sin{\gamma}\\ 0&\sin{\gamma}&\cos{\gamma}\end{pmatrix}$.
We can denote $\mathbf{A}=\mathcal{A}(\lambda_{1},\lambda_{2},\lambda_{3},\alpha,\beta,\gamma)$ for the 3D scenario.
7.1. Cage-based Deformation
First, we present the results of cage-based deformation using anisotropic Green coordinates with various control matrices. When $\mathbf{A}=\mathbf{I}$, our results align with those of the traditional isotropic technique (Lipman et al., 2008). The isotropic method ensures strict angle preservation in 2D, whereas anisotropy introduces some degree of shearing. However, we will demonstrate that this capability holds considerable practical value. We start with several representative examples to demonstrate the effects generated by different matrices $\mathbf{A}$, highlighting the diversity of deformation styles and options offered by our method. Additionally, we demonstrate scenarios in which anisotropic deformation potentials yield superior results,characterized by enhanced visual quality, tighter adherence to the deformed cage, or reduced isotropic or area distortion. These improvements are particularly evident in inputs with a high aspect ratio. Please refer to Appendix C.2 for the formula used to compute distortion. Subsequently, we will compare our method with mean value coordinates (Floater, 2003) and harmonic coordinates (Joshi et al., 2007).
Figure 4. Cage-based deformation of anisotropic Green coordinate in 2D. Different rotation matrices lead to varying “stretching” effects along different directions.
We first conduct experiments on 2D images composed of cubic and circular shapes. In Fig. 4, we present the deformation results of a grid-like image. The eigenvalues of the matrix are $1.0$ and $4.0$, while the rotation angles are $\pi/6$ and $\pi/3$. Detailed parameter settings can be found in the table of Appendix C.1. The schematic coordinate frames associated with $\lambda_{1}$, $\lambda_{2}$ and $\theta$ are also indicated in the figure. We observe that different rotation matrices lead to varying degrees of “stretching” along different directions, which intuitively demonstrates the effects of anisotropy. Fig. 5 illustrates another example involving basic shapes such as cubes and circles. The eigenvalues of the non-diagonal matrix are typically $1.0$ and $2.0$. While isotropic Green coordinates are angle-preserving and map infinitesimal circles to infinitesimal circles, their deformation effect is fixed, which limits its flexibility and diversity. In contrast, our method enables a wide variety of deformation styles. Furthermore, regarding area distortion, the leftmost image in the middle row yields the lowest value (0.420) compared to the isotropic result (0.595).
Figure 5. Cage-based deformation results of an image containing circles and cubes. Anisotropic Green coordinates brings diverse deformation effects.
Figure 6. Cage-based deformation results of a fish, an elephant, and a giraffe model, respectively.
Figure 7. Cage-based deformation results of a crab model. The tuple $(a,b)$ represents the isometric distortion and area distortion, respectively.
Figure 8. Comparison of cage-based deformation results using our method, MVC (Floater, 2003), and HC (Joshi et al., 2007).
Figure 9. Cage-based deformation results of a bar, a bench and a skybox model in 3D, respectively.
Fig. 6 shows the cage-based deformation results of a fish, an elephant, and a giraffe model, respectively. It can be observed that different matrices control the “fatness” or “thinness” of the deformed objects, thereby affording artists and animators a more versatile range of deformation choices. The crab example in Fig. 7 further demonstrates that appropriate anisotropic settings can improve deformation effects. In the figure, the tuple $(a,b)$ represents the isometric distortion and area distortion, respectively. We can observe that the anisotropic results yield more natural deformations and exhibit lower isometric and area distortion.
To further demonstrate the advantages of our method, we compare our method with mean value coordinates (MVC) (Floater, 2003) and harmonic coordinates (HC) (Joshi et al., 2007) in Fig. 8. It is evident that MVC and HC lead to noticeable shearing and unnatural deformations. Our method yields different deformation results compared to the isometric Green coordinates. Meanwhile, the quasi-conformal property of our method also guarantees a certain degree of shape preservation.
Figure 10. Results of variational shape deformation in 2D. The red dots indicate the user-specified pre-deformation positions, while the blue dots mark the target positions.
Figure 11. Results of variational shape deformation in 3D. Red and blue dots mark the spatial constraints.
We then present the results of 3D deformation. In Fig. 9, we show the results of a bar, a bench, and a skybox model, respectively. We also mark the spatial coordinate axes in the figure. In the bar model, the deformation is mainly observed along the y-axis, while the long side is oriented along the z-axis. We show the anisotropic deformation results along the y-axis or z-axis in the coefficient matrix. It can be observed that different matrices produce deformation effects with varying degrees of sharpness. In the bench model, we observe that different anisotropic matrices control the degree of upward bending in the chair’s crossbeam. In the skybox model, the first row displays results obtained using diagonal matrices, while the second row presents outcomes for non-diagonal matrices. Specifically, from left to right, the matrices are $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\pi/4,0,0)$, $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\pi/4,0,\pi/4)$ and $\mathbf{A}=\mathcal{A}(1.0,0.25,1.0,\frac{\pi}{3},\frac{\pi}{4},\frac{\pi}{6})$. These experimental results further demonstrate the wide range of deformation effects achieved by our approach.
7.2. Variational Shape Deformation
In this subsection, we conduct experiments on variational shape deformation proposed in Section 6. Specifically, we establish the deformation mapping based on anisotropic Green coordinates as Eq. (47), and optimize $\mathbf{a}$ and $\mathbf{b}$ by enforcing spatial constraints and as-rigid-as-possible constraints, as formulated in Eq. (55). The spatial constraints dictate the deformation of $\mathbf{q}_{i}$ to $\mathbf{f}_{i}$ for i=1,2,,ri=1,2,...,r. The as-rigid-as-possible constraints aim to preserve rigidity during deformation as much as possible for some selected positions. Similar to previous methods, we sample a set of points $\mathbf{m}_{i}$ for i=1,2,,di=1,2,...,d along the medial axis. The Hessian energy term and the regularization term ensure the smoothness of the deformation and the stability of optimization, respectively. Since harmonic functions satisfy the extremum principle at the boundary, this property is inherited by our approach, in the sense that boundary points map to boundary points under linear transformation. Consequently, we sample the boundary points $\mathbf{w}_{i}(i=1,2,...,k)$ near the cage faces.
Figures 10 and 11 illustrate the 2D and 3D results, respectively, which are obtained using the same input and keeping all parameters constant except for the anisotropic matrix. Red dots denote the user-specified pre-deformation positions, while blue dots mark the target positions. The 2D input cages are illustrated in orange lines. The parameters are configured as $\lambda_{1}=100,\lambda_{2}=10$ and $\lambda_{3}=0.1$. In the first two examples, we sample ten boundary points near each cage face to compute the Hessians, whereas for the final example, we sample two points near each face. The results demonstrate that varying the anisotropic matrices yields different deformation outcomes, thereby expanding the range of achievable deformation effects. In the 3D scenario, we also use red and blue points to mark the points of spatial constraints. Furthermore, we manually annotate the constraints that are occluded during rendering. The initial cages are dense, subdivided rectangular triangle meshes enclosing the entire input. We sample two points near each face to compute the Hessians. The parameters are set to $\lambda_{1}=1000,\lambda_{2}=0.5$ and $\lambda_{3}=0.001$ for the bar model (left), and to $\lambda_{1}=100,\lambda_{2}=0.05$ and $\lambda_{3}=0.001$ for the Botijo model (right). These results further demonstrate the variety of deformation effects achieved by our method.
8. Conclusion
In this work, we introduce anisotropic Green coordinates, which incorporate a symmetric positive definite coefficient matrix into the divergence operator of the Laplacian equation. This approach enables a variety of deformation styles for both cage-based deformation and variational shape deformation. We derive the Green’s identity based on the anisotropic Laplacian equation and demonstrate that our method satisfies both linear precision and translation invariance. Furthermore, anisotropic Green coordinates, along with their gradients and Hessians, admit closed-form expressions. This facilitates the efficient computation of deformation mappings and supports variational shape deformation using as-rigid-as-possible energy, while accommodating user-specified positional constraints.
The primary limitation of our method is that the matrix $\mathbf{A}$ remains globally invariant and is not optimized based on deformation constraints, which we leave to future work. Additionally, Green coordinates cannot satisfy both Dirichlet and Neumann boundary conditions simultaneously. We adopt Green coordinates as the foundation of our approach because of their theoretical elegance and strong compatibility with the concept of anisotropy. We anticipate that this framework will offer novel insights for anisotropic space deformation and inspire further research in this field.